OpenStudy (anonymous):

the first side of a triangle is 8 m shorter than the second side. the third side is 4 times as long as long as the first side. the perimeter is 26 m. find the length of each side.

OpenStudy (anonymous):

x + x-8 + 4(x-8) = 26 6x - 40 = 26 x = 11 First Side =(x-8)= 3m Second Side = x = 11 m Third Side = 4(x-8) = 12m